The question is asking if the two lines are parallel lines, coincident lines, intersecting lines, or none of those.

(1 point) The lines 4x−2y=−74x−2y=−7 and −2x+y=−2−2x+y=−2

are

a. parallel
b. coincident
c. intersecting
d. none of the above
Answer #1

The below image provides the solution.

Answer: C

- To find the lines whether, they are intersectine In Coincident or parallel Collision required is a Suppose + the equation o

2 y 4.-2y = -42-24 5-7. -2uty = -2 -2nty = -2 -2natu 2 - 2 1.бл — 2 — — 794 - +7 = 0 — 0 25 - 2-1-29 + 4 = 0 — 2 го Киеву - 2x – у +7 4 - 2 - (1) — 727 -39 +7= 0 (IT). Now, = 1therefore, i The lines are futersedia fo